Circular No. 6504

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X1730–333 R. Remillard, Massachusetts Institute of Technology (MIT), reports for

the RXTE ASM Team at MIT and the Goddard Space Flight Center: “The rapid burster X1730–333 is again seen to be in x-ray outburst. Since the end of the previous outburst (1996 May 7), the daily average measurements with the RXTE All Sky Monitor were generally upper limits near 10 mCrab. On Oct. 30, the average flux suddenly increased to 178 ± 8 mCrab (2–12 keV). During the interval Oct. 31–Nov. 7, the ASM coverage is very sparse, due to occasions of source confusion and also to the details of the RXTE pointing program. Yet there are additional detections, leading to a measurement of 115 ± 10 mCrab on Nov. 7.”

XTE J1856+053 F. E. Marshall, Laboratory for High Energy Astrophysics, Goddard

Space Flight Center (LHEA/GSFC); K. Ebisawa, Universities Space Research Association and LHEA/GSFC; R. Remillard, Center for Space Research and Department of Physics, Massachusetts Institute of Technology; and A. Valinia, National Research Council and LHEA/GSFC, report the discovery of a new x-ray source during a survey of the galactic ridge using the Proportional Counter Array on the Rossi X-ray Timing Explorer: “The 2- to 10-keV flux was ∼ 66 mCrab on each of five scans on Sept. 17 and 18. The best-fit position is α = 18^{h}56^{m}39^{s}, δ = +5^{o}19′.8 (equinox 2000.0; systematic uncertainty ∼ 0^{o}.03). Retrospective analysis of daily averaged fluxes with the ASM on RXTE shows that the source was first detected on Sept. 10, peaked during Sept. 15–17 at ∼ 35 mCrab (2–12 keV), and gradually dimmed to ∼ 7 mCrab on Oct. 31. The ASM light curve

is available at http://space.mit.edu/XTE/ASM lc.html. Observations

with imaging x-ray instruments and at other wavelengths are urged to identify this source.”

COMET 118P/SHOEMAKER-LEVY 4 Total visual magnitude estimates: Oct. 14.08 UT, 14.0 (K. Sárneczky,

Ráktanya, Hungary, 0.44-m reflector); Nov. 7.37, 13.0 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 9.15, 13.5 (Sárneczky); 11.36, 13.1 (Hale).

1996 November 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green